OpenStudy (anonymous):

Identify the center and radius for the following equation: (x+4)²+(y-7)²-4

OpenStudy (campbell_st):

rewrite the equation \[(x + 4)^2 + (y - 7)^2 = 4\] the general form of a circle is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius. hope this helps... to identify the centre and radius.

OpenStudy (anonymous):

@campbell_st thank you!! so is (x+4)^2+(y−7)^2=4 the final answer? or are there more steps after that

OpenStudy (campbell_st):

no just identify the values for h, k and r... this is a nice neat question where its written in factorised form sometimes you may need to complete the square for x and y to find the centre and radius.

OpenStudy (anonymous):

so what would the center & radius be in that case? @campbell_st

OpenStudy (campbell_st):

well what is the value of h= when you look at your equation... ?

OpenStudy (anonymous):

oh wait i think i get it now

OpenStudy (anonymous):

so they both equal 4?

OpenStudy (anonymous):

no the radius would be 2 since its squared?

OpenStudy (campbell_st):

ok... you are comparing -h = + 4 and -k = - 7 what are the values of h and k..?

OpenStudy (campbell_st):

the radius value is correct...

OpenStudy (anonymous):

what would the center be then? if -h=4?

OpenStudy (anonymous):

i don't know what positive h would be

OpenStudy (campbell_st):

-h = 4 so what is h... just multiply both sides by -1

OpenStudy (anonymous):

ohh thats right haha okay that was dumb so h=-4

OpenStudy (anonymous):

thank you so much

OpenStudy (campbell_st):

ok.. now what is k...?

OpenStudy (anonymous):

7

OpenStudy (campbell_st):

correct so the centre is (-4, 7) and radius = 2

OpenStudy (anonymous):

okay so for this problem: (x+1)²+(y+3)²=49 the center would be (1,-3) and the radius would be 7.. correct? thats what i got

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

remember -h = 1 so h = -1 centre (-1, -3) and radius is 7... good job

OpenStudy (anonymous):

thanks so much

OpenStudy (campbell_st):

glad to help

OpenStudy (anonymous):

can you possibly help me with another question similar to this? im not sure how to do it when the question is in a different format, ill post it now

OpenStudy (campbell_st):

ok... put it up